# Temporal Variation of the Coronal Parameter in a Jetted Tidal Disruption Event: Swift J1644+57

### 10 Mar 2026 v1

@@ Abstract
Tidal Disruption Events are exotic astrophysical phenomena where matter from a star or the interstellar medium is captured by a supermassive black hole. The process liberates enormous energy, within a few months to a year timescale, enough to detect dormant black holes in near as well as the farthest galaxies. We revisit the long-term spectral variabilities associated with the jetted Tidal Disruption Event Swift J1644+57 by exploring the archival X-ray data obtained with MAXI/GSC, Swift/XRT, and XMM-Newton observatories. Our analysis reveals that the spectral indices decrease non-monotonically as Swift J1644+57 evolves with time. We also find that the soft (0.3-1.5 keV) and hard (1.5-10 keV) X-ray photon counts are highly correlated with a maximum correlation coefficient of 0.95 and peak at zero lag. Moreover, the soft and hard band variabilities obtained from XMM-Newton observations are highly correlated with a Pearson cross-correlation coefficient of 0.96. This indicates that the soft and hard X-ray photons are emitted from the same site, which is most likely a Compton cloud, i.e., the corona. Assuming the hard X-ray photons originate from the corona, we find that the coronal parameter undergoes rapid expansion during the early phases when accompanied by a relativistic jet launching and subsequently evolves toward a state of saturation with minor fluctuations in the latter stages. The temporal variation of the coronal size is consistent with a simple theoretical conjecture. We also discuss the application of our analytical outcomes to other jetted and non-jetted tidal disruption events.
@@ I Introduction
There are differences in X-ray spectral hardness between jetted TDEs and non-jetted TDEs. Non-jetted TDEs often exhibit softer spectra, where photon index $\Gamma>2.5$ is associated with a thermal component (Komossa, 2017). Contrary to that, jetted TDEs have much harder spectra with $\Gamma\leq 2$. In addition, no clear evidence of a thermal component could be observed from the X-ray spectral energy distribution (Mangano et al., 2016; Seifina et al., 2017). The domain of jetted TDEs began in March 2011 with the discovery of Swift J1644+57. The origin of high-energy $\gamma$-rays is traced back to a galaxy at $z=0.35$ (Levan et al., 2011). The peak X-ray luminosity exceeded the Eddington limit with the inferred black hole mass $3\times 10^{6}M_{\odot}$ (Bloom et al., 2011b; Burrows et al., 2011; Levan et al., 2011). The source was detected in radio (Cendes et al., 2014; Berger et al., 2012; Zauderer et al., 2013), infra-red (Wiersema et al. 2012); X-rays (Reis et al., 2012; Saxton et al., 2012; González-Rodríguez et al., 2014; Mangano et al., 2016; Zauderer et al., 2013), $\gamma$-rays (Bloom et al., 2011b; Burrows et al., 2011). Unlike X-ray luminosity, which decayed over the next 500 days, the radio luminosity increased (Berger et al., 2012; Zauderer et al., 2013)).
@@
A quasi-periodic oscillation (QPO) of around 200 s within the first 20 days from the peak was discovered from the X-ray light curves in the 2-10 keV energy band from Suzaku and XMM-Newton observations (Reis et al., 2012; Saxton et al., 2012). Similar QPOs in X-rays have been observed for two other TDE candidates, namely 2XMM J123103.2+110648 (Lin et al., 2013) and ASASSN-14li (Pasham et al., 2019). Seifina et al. (2017) showed that the spectral hardness of the source evolved from $\Gamma\sim 1.8$ to $\Gamma\sim 1.2$ for the first 200 days since the outburst. However, the variation of spectral index met sudden variations associated with the dips found in the light curve (Mangano et al., 2016). From the X-ray spectral shape, luminosity, and the absence of the thermal component, Zauderer et al. (2013) asserted that the X-ray photons of Swift J1644+57 could not have originated from the accretion disk. Rather, the origin of the X-rays could be from the internal dissipation within the jet at a large distance ($r\sim 10^{4}-10^{5}r_{\rm g}$) from the SMBH. Contrary to their idea, Lei et al. (2016) suggested the X-rays from the off-axis jetted-TDE candidate IGR J12580+0134 is the disk-corona origin. Stronger evidence of the X-ray origin of Swift J1644+57 came from Kara et al. (2016), where reverberation lag was observed. Similar to the QPO detection of Swift J1644+57, Nikołajuk and Walter (2013) suggested detection of QPO within 0.00080.0008-$0.004$  Hz, which is consistent with a black hole of mass $\sim 10^{5}M_{\odot}$.
@@
There are many arguments that the QPOs can bridge the mass gap between the accretion processes of galactic, stellar-mass black holes, and SMBHs (Vaughan and Uttley, 2005). However, very few of them are detected with statistical confidence for AGNs (Gierliński et al., 2008; Alston et al., 2015; Ashton and Middleton, 2021). There are many challenges for observing QPOs originating from AGNs, apart from achieving a statistically significant signal to noise ratio (Vaughan and Uttley, 2006). Considering the mass of Swift J1644+57, the 200 s QPO found in the X-ray light curve requires a scale of $\mathcal{O}(10)\,r_{\rm S}$ for a non-rotating SMBH. This scale fits within the range of the typical coronal size of TDEs as proposed by Mummery and Balbus (2021), and is also consistent with the corona size, which is observationally suggested for many AGNs (Alston et al., 2020). Thus, we expect to find a reasonable explanation of hard X-ray variations through the disk-corona geometry by analyzing the existing data of jetted TDEs. The most promising mechanism for the luminous emission of TDEs is a thermal blackbody emission. Gezari (2021) suggested that there is a bimodal distribution of the TDE thermal emission, indicating that the soft X-ray photons are produced from the inner disk region, whereas the optical/UV photons are emitted from the outflow’s photosphere. The thermal photons produced from the disk can be a good candidate source for the Compton scattering. This suggests a disk-corona system can exist in X-ray TDEs.
@@ II Data Analysis Process
We used the archival data of XMM-Newton and Swift/XRT obtained through HEASARC [^1] http://heasarc.gsfc.nasa.gov/. We reprocessed all data using HEAsoft v6.26.1 (Arnaud, 1996), which includes XSPEC v12.10.1f. We use the following cosmological parameters in this work: $H_{0}$ = 70 km s-1 Mpc-1, $\Lambda_{0}$ = 0.73, $\Omega_{M}$ = 0.27 (Bennett et al., 2003). With the cosmological parameters, the luminosity distance of Swift J1644+57 was considered 1408 Mpc (Levan et al., 2011).
@@ III.2.1 Optxagnf
The Optxagnf model, proposed by Done et al. (2012), provides an in-depth understanding of the spectral parameters as the parameters are generated from a physical picture of the accretion disk around the AGNs. The model-fitted parameters include the luminosity, coronal size, spin of the central black hole, soft-excess temperature, optical depth, spectral index, and the fraction of Comptonized photons. It should be noted that Optxagnf is the only publicly available model to fit the X-ray spectra of AGNs where coronal size could be estimated. We employed this model to fit the XRT spectra. The model in xspec reads as: TBabs*zTBabs*(optxagnf).
@@
The model spectrum computes both soft-excess and Comptonized photons. The accretion rate governed by the mass of the central black hole determines the luminosity of the spectra in units of Eddington luminosity. The geometry of the Optxagnf model separates the accretion disk into three regions. First, the disk emission is produced through a color temperature-corrected blackbody emission from radii $R_{\rm out}>r>R_{\rm cor}$, where $R_{\rm out}$ and $R_{\rm cor}$ are the outer edge of the disk and the corona, respectively. Second, at $r<R_{\rm cor}$, the emission is dominated by Comptonized emission from a warm and optically thick medium. The third region where the inverse Comptonization occurs is in the optically thin corona located around the disk, and the fraction of inverse Comptonization could be estimated from the $f_{pl}$ parameter while fitting the spectra. The temperature and optical depth of the warm, optically thick medium could be found from $kT_{s}$ and $\tau$, respectively. The model calculates the primary emission of Comptonized photons using the Nthcomp model, where the seed photon temperature is kept fixed at $r=R_{\rm cor}$, and the electron temperature of the optically thin hot corona is frozen at 100 keV. We freeze the mass of the black hole at $3\times 10^{6}~M_{\odot}$ (Bloom et al., 2011b). The galactic $N_{H}$ along the line of Swift J1644+57 remained frozen while performing the fitting. According to the model, it was suggested to keep the normalization frozen at unity. The redshift $z=0.354$ and distance $d=1402.8$ Mpc are also kept constant (Levan et al., 2011). From Tab 3, we can see that the power law model fits the spectra of Swift J1644+57 well, and the soft X-ray excess is absent in the spectra. Thus, we freeze the $kT_{s}$ at 0.1 keV and $\tau$ at a maximum value of 5.0. Initially, we left the $f_{pl}$ thawed. We observed that the $f_{pl}$ values saturated for most of the cases. Later, better fits were obtained while fitting the spectra with variable $kT_{s}$ and $\tau$. We observed that the warm corona temperature and optical depth remained high for all of the observations except XRT3. We found that the $kT_{s}$ gradually increases as the event progresses and remains within $0.59\leq kT_{s}\leq 5.73$. On the other hand, the variation of optical depth remained non-monotonic within the limit $1.25\leq\tau\leq 7.83$. The variation of $f_{pl}$ remained within the range of $0.48\leq f_{pl}\leq 0.93$ during our observation period. For TDEs, this is expected and is conjectured by Chakraborty et al. (2026). Optxagnf fitted spectra for three epochs are presented in Fig. 3, and the variation of fitted parameters with time is plotted in the right panel of Fig. 5 and are presented in the Table 4.
@@
We fitted the 0.3–10 keV stacked Swift spectra for the entire episode of TDE. The spectra fitted alone with the Optxagnf model provided good statistics with $\chi^{2}_{red}\sim 1$. The model-fitted spectral parameters are presented in Table 4. We found the luminosity of the source exceeded nearly 500 times the Eddington luminosity during the initial few days of the event. This agrees with the earlier studies like Zauderer et al. (2013) and is also seen from the current fitting with Optxagnf model. Over time, the luminosity of the source decreased nearly 200 times from the initial phase, but never went into the sub-Eddington regime during the period of our observation. The variation of $N_{H}$ remained similar to what has been observed during the fitting with the powerlaw model. The spin of the central black hole remained within the range of 0.4<a<0.80.4<a*<0.8, making it an intermediate spinning black hole. During the initial days, the coronal parameter of the source was observed at $31\pm 2~r_{\rm g}$. The right panel of Fig. 4 represents the confidence contours of spin ($a$) vs coronal radius ($r_{cor}$). The coronal radius was found to vary within $31<R_{cor}<65~r_{\rm g}$. We only used the stacked Swift data for Optxagnf fitting as it provides better data quality to understand long-timescale variations of Swift J1644+57 during its tidal disruption episode.
@@ IV Discussion
We studied Swift J1644+57 during its tidal disruption event of 2011-2012 using Swift and XMM-Newton data within the energy range 0.3-10 keV. We employed the powerlaw as our phenomenological model and a physical model, such as Optxagnf , to understand the behavior of the source as the event progresses. Being a jetted-TDE, the source exhibited variability in both the spectral and temporal domains and is one of the most studied tidal disruption events to date. During our analysis, we found that there are several similarities between TDEs and the declining phase of an outbursting GBH. The similarities between galactic black hole (GBH) transients and TDEs are seen from their spectral energy distributions and the decaying luminosity with time. The X-ray spectral indices of GBHs are found to decrease in the luminosity decay phase (McClintock et al., 2009). Likewise, the decreasing tendency of the spectral index has also been observed in Swift J1644+57 (Seifina et al., 2017). Recent NICER observations (Gendreau et al., 2012) of MAXI J1820+070 provided evidence of a dynamically contracting corona for the onset stage of the flare (Kara et al., 2019). Jana et al. (2016) explained, using the two-component advective flow model (Chakrabarti and Titarchuk, 1995), why the Compton cloud expanded at the declining phase of a low inclination GBH, MAXI J1836-194, during the 2011 outburst. The variation of the corona size is also observed for AGN IRAS 13224-3809 (Alston et al., 2020). Apart from coronal size variation, relativistic jet launching is another phenomenon that GBH transients and TDEs have in common (Mirabel and Rodríguez, 1994; Patra et al., 2019; Espinasse et al., 2020; Blandford et al., 2019; Zauderer et al., 2011).

### 07 Apr 2026 v2

# Temporal Variation of the Coronal Radius Parameter in a Jetted Tidal Disruption Event: Swift J1644+57

@@ Abstract
Tidal Disruption Events are exotic astrophysical phenomena where matter from a star or the interstellar medium is captured by a supermassive black hole. The process liberates enormous energy, within a few months to a year timescale, enough to detect dormant black holes in near as well as the farthest galaxies. We revisit the long-term spectral variabilities associated with the jetted Tidal Disruption Event Swift J1644+57 by exploring the archival X-ray data obtained with Swift/XRT and XMM-Newton observatories. Our analysis reveals that the spectral indices decrease non-monotonically as Swift J1644+57 evolves with time. We also find that the soft (0.3-1.5 keV) and hard (1.5-10 keV) X-ray photon counts are highly correlated with a maximum correlation coefficient of 0.95 and peak at zero lag. Moreover, the soft and hard band variabilities obtained from XMM-Newton observations are highly correlated with a Pearson cross-correlation coefficient of 0.96. This indicates that the soft and hard X-ray photons are emitted from the same site, which is most likely a Compton cloud, i.e., the corona. Assuming the hard X-ray photons originate from the corona, we find that the coronal parameter undergoes rapid expansion during the early phases when accompanied by a relativistic jet launching and subsequently evolves toward a state of saturation with minor fluctuations in the latter stages. The temporal variation of the coronal radius parameter ($R_{cor}$) is consistent with a simple theoretical conjecture. We also discuss the application of our analytical outcomes to other jetted and non-jetted tidal disruption events.
@@ I Introduction
There are differences in X-ray spectral hardness between jetted TDEs and non-jetted TDEs. Non-jetted TDEs often exhibit softer spectra, where photon index $\Gamma>2.5$ is associated with a thermal component (Komossa, 2017). Contrary to that, jetted TDEs have much harder spectra with $\Gamma\leq 2$. In addition, no clear evidence of a thermal component could be observed from the X-ray spectral energy distribution (Mangano et al., 2016; Seifina et al., 2017). The domain of jetted TDEs began in March 2011 with the discovery of Swift J1644+57. The origin of high-energy $\gamma$-rays is traced back to a galaxy at $z=0.35$ (Levan et al., 2011). The peak X-ray luminosity exceeded the Eddington limit with the inferred black hole mass $3\times 10^{6}M_{\odot}$ (Bloom et al., 2011b; Burrows et al., 2011; Levan et al., 2011). The source was detected in radio (Cendes et al., 2014; Berger et al., 2012; Zauderer et al., 2013), infra-red (Wiersema et al. 2012); X-rays (Reis et al., 2012; Saxton et al., 2012; González-Rodríguez et al., 2014; Mangano et al., 2016; Zauderer et al., 2013), $\gamma$-rays (Bloom et al., 2011b; Burrows et al., 2011). Unlike X-ray luminosity, which decayed over the next 500 days, the radio luminosity increased (Berger et al., 2012; Zauderer et al., 2013) as the event progressed.
@@
Seifina et al. (2017) showed that the spectral hardness of the source evolved from $\Gamma\sim 1.8$ to $\Gamma\sim 1.2$ for the first 200 days since the outburst. However, the variation of spectral index met sudden fluctuations associated with the dips found in the light curve (Mangano et al., 2016). From the X-ray spectral shape, luminosity, and the absence of the thermal component, Zauderer et al. (2013) asserted that the X-ray photons of Swift J1644+57 could not have originated from the accretion disk. Rather, the origin of the X-rays could be from the internal dissipation within the jet at a large distance ($r\sim 10^{4}-10^{5}r_{\rm g}$) from the SMBH. Contrary to their idea, Lei et al. (2016) suggested the X-ray emission from the off-axis jetted-TDE candidate IGR J12580+0134 originates from the disk-corona. Beamed accretion, much like in the case of Blazers, could stand out as a possible scenario for the hyper accretion rate. However, in the case of Swift J1644+57, the luminosity kept falling with decreasing X-ray photon index ($\Gamma$). This observed correlation between luminosity and $\Gamma$ opposes the “harder when brighter” (Krawczynski et al., 2004; Gliozzi et al., 2006; Zhang et al., 2006; Pandey et al., 2017) nature of the X-ray bright Blazars and aligns with the accretion disk dominated Seyferts or galactic black holes (GBHs) (e.g., Yang et al., 2015; Jana, 2022; Jana et al., 2026). Stronger evidence of the accretion disk origin of the X-ray emission in Swift J1644+57 came from Kara et al. (2016), where reverberation lag was observed.
@@
A quasi-periodic oscillation (QPO) of 200 s, within the first 20 days from the peak luminosity, was discovered from the X-ray light curves in the 2-10 keV energy band, using the Suzaku and XMM-Newton observations (Reis et al., 2012; Saxton et al., 2012). Similar QPOs in X-rays have been observed for two other TDE candidates, namely 2XMM J123103.2+110648 (Lin et al., 2013) and ASASSN-14li (Pasham et al., 2019). Nikołajuk and Walter (2013) suggested the detection of a QPO within 0.00080.0008-$0.004$  Hz for IGR J12580+0134, where the timescale is consistent for a TDE with central black hole mass $\sim 10^{5}M_{\odot}$. There are many arguments that the QPOs can bridge the mass gap between the accretion processes of galactic, stellar-mass black holes, and SMBHs (Vaughan and Uttley, 2005). However, very few of them are detected with statistical confidence for AGNs (Gierliński et al., 2008; Alston et al., 2015; Ashton and Middleton, 2021). There are many challenges for observing QPOs originating from AGNs, apart from achieving a statistically significant signal to noise ratio (Vaughan and Uttley, 2006). Considering the mass of Swift J1644+57, the 200 s QPO found in the X-ray light curve requires a scale of $\mathcal{O}(10)\,r_{\rm s}$ for a non-rotating SMBH. This scale fits within the range of the typical coronal radius parameter (hereafter $R_{cor}$) of TDEs as proposed by Mummery and Balbus (2021), and is also consistent with the corona size, which is observationally suggested for many AGNs (Alston et al., 2020). Thus, we expect to find a reasonable explanation of hard X-ray variations through the disk-corona geometry by analyzing the existing data of jetted TDEs. The most promising mechanism for the luminous emission of TDEs is a thermal blackbody emission. Gezari (2021) suggested that there is a bimodal distribution of the TDE thermal emission, indicating that the soft X-ray photons are produced from the inner disk region, whereas the optical/UV photons could be emitted from the photospheric outflows. The thermal photons produced from the disk can be a good candidate source for the Compton scattering. This suggests a disk-corona system can exist in X-ray TDEs.
@@ II Data Analysis Process
We used the archival data of XMM-Newton and Swift/XRT obtained through HEASARC [^1] http://heasarc.gsfc.nasa.gov/. We reprocessed all data using HEAsoft v6.26.1 (Arnaud, 1996), which includes XSPEC v12.10.1f. We use the following cosmological parameters in this work: $H_{0}$ = 70 km s-1 Mpc-1, $\Lambda_{0}$ = 0.73, $\Omega_{M}$ = 0.27 (Bennett et al., 2003). With the cosmological parameters, the luminosity distance of Swift J1644+57 was considered 1402.8 Mpc (Levan et al., 2011).
@@ III.2.1 Optxagnf
The Optxagnf model, proposed by Done et al. (2012), provides an in-depth understanding of the spectral parameters as the parameters are generated from a physical picture of the accretion disk around the AGNs. The model-fitted parameters include the luminosity, coronal radius parameter ($R_{cor}$), spin of the central black hole, soft-excess temperature, optical depth, spectral index, and the fraction of Comptonized photons. It should be noted that Optxagnf is the only publicly available model to fit the X-ray spectra of AGNs where $R_{cor}$ could be estimated. We employed this model to fit the XRT spectra. The model in xspec reads as: TBabs*zTBabs*(optxagnf).
@@
The model spectrum computes both soft-excess and Comptonized photons. The accretion rate governed by the mass of the central black hole determines the luminosity of the source (in units of Eddington luminosity - $log(L/L_{Edd})$). The geometry of the Optxagnf  model separates the accretion disk into three regions. First, the disk emission is produced through a color temperature-corrected blackbody emission from radii $R_{\rm out}>r>R_{\rm cor}$, where $R_{\rm out}$ and $R_{\rm cor}$ are the outer edge of the disk and the corona, respectively. Luminosity is computed keeping the distance and mass constant, while the mass-to-energy conversion efficiency factor ($\eta$) depends on the spin. Second, at $r<R_{\rm cor}$, the emission is dominated by Comptonized emission from a warm and optically thick medium. The third region where inverse Comptonization occurs is the optically thin corona around the disk, and the fraction of inverse Comptonization can be estimated from the $f_{pl}$ parameter during the spectral fitting. The temperature and optical depth of the warm, optically thick medium could be found from $kT_{s}$ and $\tau$, respectively. The model calculates the primary emission of Comptonized photons using the Nthcomp model, where the seed photon temperature is kept fixed at $r=R_{\rm cor}$, and the electron temperature of the optically thin hot corona is frozen at 100 keV. We freeze the mass of the black hole at $3\times 10^{6}~M_{\odot}$ (Bloom et al., 2011b). The Galactic $N_{H}$ along the line of Swift J1644+57  remained frozen at $0.016\times 10^{22}\mathrm{cm}^{-2}$ while performing the fitting. According to the model, it was suggested to keep the normalization frozen at unity. The redshift $z=0.354$ and distance $d=1402.8$ Mpc are also kept fixed (Levan et al., 2011). The $log(r_{out})$ is kept fixed at 5 since the apogee of extremely elongated TDEs, capable of explaining lengthy disruption events (Hayasaki et al., 2013; Dai et al., 2015; Hayasaki et al., 2018), could reach up to $10^{4}-10^{5}\,r_{\rm g}$. The remaining parameters are left free to vary. It should be noted that the reported luminosities in Table 4 are isotropic and model-dependent without employing the beaming corrections. From Table 3, we can see that the power law model fits the spectra of Swift J1644+57  well, and the soft X-ray excess is absent in the spectra. Thus, we freeze the $kT_{s}$ at 0.1 keV and $\tau$ at a maximum value of 5.0. Initially, we left the $f_{pl}$ thawed. We observed that the $f_{pl}$ values saturated in most cases. Later, better fits were obtained while fitting the spectra with variable $kT_{s}$ and $\tau$ and is presented in Table 4. We observed that the warm corona temperature and optical depth remained high across all observations except XRT3. We found that the $kT_{s}$ gradually increases as the event progresses and remains within $0.59\leq kT_{s}\leq 5.73$. On the other hand, the variation of optical depth remained non-monotonic within the limit $1.25\leq\tau\leq 7.83$. The variation of $f_{pl}$ remained within the range of $0.48\leq f_{pl}\leq 0.93$ during our observation period. For TDEs, this is expected and is conjectured by Chakraborty et al. (2026). Optxagnf fitted spectra for three epochs are presented in Fig. 3, and the variation of fitted parameters with time is plotted in the right panel of Fig. 5 and are presented in the Table 4.
@@
We fitted the 0.3–10 keV stacked Swift spectra for the entire episode of TDE. The spectra fitted alone with the Optxagnf model provided good statistics with $\chi^{2}_{red}\sim 1$. The model-fitted spectral parameters are presented in Table 4. We found the luminosity of the source exceeded nearly 500 times the Eddington luminosity during the initial few days of the event. This agrees with the earlier studies like Zauderer et al. (2013) and is also seen from the current fitting with Optxagnf model. Over time, the luminosity of the source decreased nearly 200 times from the initial phase, but never went into the sub-Eddington regime during the period of our observation. The variation of $N_{H}$ remained similar to what has been observed during the fitting with the powerlaw model. The spin of the central black hole remained within the range of $0.4<a^{*}<0.8$, making it an intermediate spinning black hole. During the initial days, the coronal parameter of the source was observed at $31\pm 2~r_{\rm g}$. The right panel of Fig. 4 represents the confidence contours of spin ($a$) vs $R_{cor}$. The coronal radius parameter was found to vary within $31<R_{cor}<65~r_{\rm g}$. We only used the stacked Swift data for Optxagnf fitting as it provides better data quality to understand the long-timescale variations of Swift J1644+57 during its tidal disruption episode.
@@ IV Discussion
We studied Swift J1644+57 during its tidal disruption event of 2011-2012 using Swift and XMM-Newton data within the energy range 0.3-10 keV. We employed the powerlaw as our phenomenological model and a physical model, such as Optxagnf, to understand the behavior of the source as the event progresses. Being a jetted-TDE, the source exhibited variability in both the spectral and temporal domains and is one of the most studied tidal disruption events to date. During our analysis, we found that there are several similarities between TDEs and the declining phase of an outbursting GBH. The similarities between GBH transients and TDEs are seen from their spectral energy distributions and the decaying luminosity with time. The X-ray spectral indices of GBHs are found to decrease in the luminosity decay phase (McClintock et al., 2009). Likewise, the decreasing tendency of the spectral index has also been observed in Swift J1644+57 (Seifina et al., 2017). Recent NICER observations (Gendreau et al., 2012) of MAXI J1820+070 provided evidence of a dynamically contracting corona for the onset stage of the flare (Kara et al., 2019). Jana et al. (2016) explained, using the two-component advective flow model (Chakrabarti and Titarchuk, 1995), why the Compton cloud expanded at the declining phase of a low inclination GBH, MAXI J1836-194, during the 2011 outburst. The variation of the corona size is also observed for AGN IRAS 13224-3809 (Alston et al., 2020). Apart from $R_{cor}$ variation, relativistic jet launching is another phenomenon that GBH transients and TDEs have in common (Mirabel and Rodríguez, 1994; Patra et al., 2019; Espinasse et al., 2020; Blandford et al., 2019; Zauderer et al., 2011).